### v1

# Attribute-driven Disentangled Representation Learning for Multimodal Recommendation

Abstract.
Recommendation algorithms forecast user preferences by correlating user and item representations derived from historical interaction patterns. In pursuit of enhanced performance, many methods focus on learning robust and independent representations by disentangling the intricate factors within interaction data across various modalities in an unsupervised manner. However, such an approach obfuscates the discernment of how specific factors (e.g., category or brand) influence the outcomes, making it challenging to regulate their effects. In response to this challenge, we introduce a novel method called Attribute-Driven Disentangled Representation Learning (short for AD-DRL), which explicitly incorporates attributes from different modalities into the disentangled representation learning process. By assigning a specific attribute to each factor in multimodal features, AD-DRL can disentangle the factors at both attribute and attribute-value levels. To obtain robust and independent representations for each factor associated with a specific attribute, we first disentangle the representations of features both within and across different modalities. Moreover, we further enhance the robustness of the representations by fusing the multimodal features of the same factor. Empirical evaluations conducted on three public real-world datasets substantiate the effectiveness of AD-DRL, as well as its interpretability and controllability.
Multimodal, Attribute, Disentangled Representation, Recommendation
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Personalizationccs: Information systems Recommender systemsccs: Information systems Collaborative filtering
1. Introduction
Recommender systems (RS) are integral to a myriad of online platforms, spanning E-commerce to advertising, facilitating users in pinpointing items aligned with their preferences. Given their pivotal role, numerous efforts have been dedicated to developing advanced recommendation models to improve their performance. Among them, Collaborative Filtering (CF) based models (Rendle et al., 2009; Rendle, 2010; He et al., 2017; Hsieh et al., 2017; Wang et al., 2019; He et al., 2020; Wang et al., 2020) have achieved great success by exploiting the user-item interaction data to learn user and item representation. However, these models can easily encounter the sparsity problem in practical scenarios due to their exclusive dependence on interaction data. To alleviate the data sparsity problem in recommendation, side information (e.g., attributes, user reviews and item images), which contains rich information associated with users or items, is often used to enhance the representation learning of users and items (Zhang et al., 2016, 2017a; Liu et al., 2019; Wei et al., 2019, 2020; Liu et al., 2022, 2023).
It is well recognized that the entangled representations of users and items are infeasible to directly capture fine-grained user preferences across diverse factors, thereby constraining both the efficacy and interpretability of recommender systems (Ma et al., 2019; Wang et al., 2020). In recent years, the study of disentangled representation learning has garnered significant attention in diverse fields, notably in computer vision (Higgins et al., 2017; Liu et al., 2020; Hsieh et al., 2018), due to its capability to identify and disentangle the underlying factors behind data. Empirical evidence suggests that disentangled representations exhibit enhanced robustness, particularly in intricate application contexts. Hence, many recommendation methods adopt disentangled representation learning techniques to learn robust and independent representations, ultimately enhancing recommendation performance (Ma et al., 2019; Mu et al., 2022; Wang et al., 2020; Tran and Lauw, 2022; Liu et al., 2022). For example, Ma et al. (Ma et al., 2019) employed disentangled representations to capture user preferences regarding different concepts associated with user intentions. Wang et al. (Wang et al., 2020) introduced a GCN-based model that produces disentangled representations by modelling a distribution over intents for each user-item interaction, exploring the diversity of user intentions on adopting items. However, these methods only concentrate on disentangling the user and item representations based on their ID embeddings. To exploit the difference between the factors behind data of various modalities, Liu et al. (Liu et al., 2022) estimated the users’ attention weight to underlying factors of different modalities, utilizing a sophisticated attention-driven module.
Despite the considerable advancements brought about by disentangling techniques in recommendation systems, existing studies often disentangle both users and items into latent factors. This methodology inherently constrains the model’s interpretability and controllability. For example, consider using disentangled representations to elucidate the diverse factors, such as style, brand, popularity, and price, that shape user preferences in dress selection. The inherent abstraction of latent factors from current disentangling methods makes it arduous to pinpoint which factor represents each specific dress attribute. In particular, one factor might be loosely related to a combination of brand and price, while another factor might represent a mixture of style and popularity. This ambiguity in the latent factors limits the interpretability of the recommendation system, making it difficult to understand why a particular item was recommended to a user. Moreover, this lack of clarity also affects the controllability of the recommendation system. Suppose a user wants to receive recommendations specifically for dresses without considering her preferences for price and popularity. It would be challenging to adjust the recommendation system to focus on those specific preferences, as the latent factors are not clearly tied to these attributes.
Item attributes manifested in various modalities can enrich the recommendation system by offering diverse and complementary information. For example, the textual data might explicitly mention the brand or price, while visual data can reveal visual attributes, such as the category of items. Additionally, the popularity of items can be derived from the statistical information of interaction data. In fact, attributes represent specific, meaningful properties or characteristics of items. Consequently, using attributes to guide the disentanglement process is a promising way to improve the interpretability and controllability of conventional multimodal recommendation methods. In this paper, we propose an Attribute-Driven Disentangled Representation Learning method (AD-DRL for short), which disentangles factors in user and item representations across various modalities at different levels of attribute granularity. To obtain robust and independent representations for each factor associated with a specific attribute, we first disentangle the representations of features within and across different modalities. This process is guided by high-level attributes (e.g., category and popularity), which help reveal the underlying relationships between factors. Following this, we further enhance the robustness of the representations by fusing the multimodal features of the same factor. This step involves exploiting the relationships among representations of the same factor by leveraging low-level attributes (e.g., category attribute values for a clothing dataset, such as jeans, jackets, and dresses), resulting in finer-grained and more comprehensive disentangled representations. To validate the effectiveness of our method, we conduct extensive experiments and ablation studies on three real-world datasets. Experimental results demonstrate the superiority of our method AD-DRL compared to existing methods and showcase its capability in terms of interpretability and controllability.
In summary, the contribution of this paper is threefold:
  • In this paper, we highlight the limitations of traditional disentangled representation learning in multimodal recommendation systems. To overcome these shortcomings, we introduce AD-DRL, which improves interpretability and controllability by employing attributes to disentangle factors in user and item representations.
  • To achieve robust and independent representations, we assign a specific attribute to each factor in multimodal features and disentangle factors at both the attribute and attribute-value levels.
  • We conduct extensive experiments on three real-world datasets to validate the effectiveness of our method. The experimental analysis demonstrates the interpretability and controllability of our model.
2. Related Work
In this section, we present a concise overview of the latest advancements in multimodal collaborative filtering techniques and disentangled representation learning in recommendation, as they hold considerable pertinence to our study.
2.1. Multimodal Collaborative Filtering
Traditional Collaborative Filtering (CF) (Sarwar et al., 2001; Ebesu et al., 2018; Rendle et al., 2009) methods primarily rely on user-item interaction behaviour to learn representations of users and items. Consequently, the recommendation performance is negatively impacted when encountering users and items with limited interactions. To mitigate the challenges posed by data sparsity (McAuley and Leskovec, 2013; Kalantidis et al., 2013; Tan et al., 2016), recent research has incorporated multimodal information into recommendation systems. The multimodal information, such as reviews and images, provides valuable information of user preference and item characteristics, which can supplement historical user-item interactions and thus enhance recommendation performance (He and McAuley, 2016; Zhang et al., 2017a; Liu et al., 2019; Wei et al., 2019, 2020; Liu et al., 2022, 2023).
Early CF models like matrix factorization (MF) (Koren et al., 2009) project the IDs of users and items to embeddings by minimizing the error of reconstructing the user-item interaction matrix. Previous studies integrate multimodal information into the MF-based method in a straightforward manner. For instance, VBPR (He and McAuley, 2016) directly concatenates the visual features learned from item images with collaborative features as the joint item representation and feeds it into the MF module. With the success of deep learning techniques in modeling complex interaction behaviors (He et al., 2017) and the relation between various multimodal features (Nie et al., 2016), a lot of deep learning techniques have been introduced into multimodal recommendation (He et al., 2017; Tay et al., 2018). For example, MAML (Liu et al., 2019) first fuses the item’s multimodal features and user features, then feeds it into an attention module to capture users’ diverse preferences. VECF (Chen et al., 2019) constructs a visually explainable collaborative filtering model using a multimodal attention network, facilitating the integrated coupling of diverse feature modalities. More recently, graph convolutional networks (GCNs) (Hamilton et al., 2017) have demonstrated their power capability of representation learning for recommendation (Wang et al., 2019; He et al., 2020; Zhou et al., 2023; Wei et al., 2023). Based on the GCN structure, MMGCN (Wei et al., 2019) constructs a user-item bipartite graph to learn representations of each modality and then fuse the modality representation together as the final representation. As there are false-positive edges exist which will hurt the performance of recommendation, GRCN (Wei et al., 2020) exploits the rich multimodal content of items to refine the structure of the interaction graph.
2.2. Disentangled Representation Learning
Disentangled representation learning, which seeks to identify and separate underlying explanatory factors embedded in data, has garnered significant interest, particularly within the realm of computer vision (Higgins et al., 2017; Liu et al., 2020; Ge et al., 2021). For instance, the beta-VAE method (Higgins et al., 2017) employs a constrained variational framework to learn disentangled representations of fundamental visual concepts. Meanwhile, IPGDN (Liu et al., 2020) automatically uncovers independent latent factors present in graph data.
Owing to the successful application of disentangled representation learning in various domains, numerous studies have concentrated on learning disentangled representations for users and items in recommendation systems in recent years (Ma et al., 2019; Tran and Lauw, 2022; Wang et al., 2020; Mu et al., 2022; Liu et al., 2022; Wang et al., 2023). Following the previous work in computer vision, the initial attempts take Variational Auto-encoder (VAE) (Kingma and Welling, 2014) to learn disentangled representations. For example, MacridVAE (Ma et al., 2019) captured user preferences regarding the different concepts associated with user intentions separately. Beyond the disentangled representations derived from user-item interaction modeling, ADDVAE (Tran and Lauw, 2022) acquires an additional set of disentangled user representations from textual content and subsequently aligns both sets of representations. For studying the diversity of user intents on adopting the items, DGCF (Wang et al., 2020) employs a graph disentangled module to iteratively refine the intent-aware interaction graph and factorial representations for recommendations. In order to model the users’ various preferences on different factors of each modality, DRML (Liu et al., 2022) estimates the user’s attention weight to underlying factors of different modalities with an attention module. As the learned disentangled representations lack clear meanings, KDR (Mu et al., 2022) harnesses the power of knowledge graphs (KGs) to guide the disentangled representation learning process, ensuring that the resulting disentangled representations are associated with semantically meaningful information extracted from the KGs.
Although the existing models achieve performance improvement by disentangled representation learning, they all face the problem that they all disentangle the user and item representation into latent factors without clarifying the semantic meaning of each factor.
3. Method
Figure 1. High-level and low-level attribute-driven disentangled representation learning module of our proposed AD-DRL. To disentangle attribute factors in multimodal features, (a) Intra-modality disentanglement module exploits the difference between attribute factors (e.g., price, brand, category and popularity) within the same modality feature, (b) Inter-modality disentanglement module utilizes the consistency of the same attribute factor in different modality features, and (c) low-level disentangled representation learning module leverages the intrinsic relationships between items sharing the same attribute value (e.g., the popularity value of different levels: Super Popular, Popular, Moderate, Emerging, and Unknown).
In this section, we first formally formulate the problem statement of multimodal recommendation and disentangled representation learning in Section 3.1. We then detail our designed attribute-driven disentangled representation learning in Section 3.2. Lastly, we describe the preference prediction process of our model in Section 3.3.
3.1. Preliminaries
Before delving into the details of our model, we first outline the problem setting and the intuition of attribute-driven disentanglement.
3.1.1. Problem Setting
Given a set of users $u\in\{1,2,\cdots,N_{u}\}$ and items $i\in\{1,2,\cdots,N_{i}\}$, we utilize two types of information to learn user and item representations: (1) A user-item interaction matrix $\mathbf{R}\in\mathbb{R}^{N_{u}\times N_{i}}$, where each entry $r_{ui}\in\mathbf{R}$ represents the implicit feedback (e.g., clicks, likes or purchases) of user $u$ to item $i$. $r_{ui}=1$ indicates that user $u$ has interacted with item $i$ and $r_{ui}=0$ indicates that there is no interaction between user $u$ and item $i$ in the observed data; and (2) multimodal features, which mainly consist of two types of features associated with items, namely reviews and images. By leveraging both user-item interactions and multimodal features, we aim to learn robust representations of users and items, thereby predicting a user’s preference to an uninteracted item.
3.1.2. Notations.
Similar to previous work (Wei et al., 2020; Liu et al., 2022), each user $u$ and item $i$ are assigned with a unique ID and respectively represented by a vector $\bm{p_{u}}\in\mathbb{R}^{d}$ and $\bm{q_{i}}\in\mathbb{R}^{d}$, which are randomly initialized in our model. For the review and image information, we use the BERT (Devlin et al., 2019) and the ViT (Dosovitskiy et al., 2021) to extract the raw textual features $\bm{F_{t}}\in\mathbb{R}^{d_{0}}$ and visual features $\bm{F_{v}}\in\mathbb{R}^{d_{0}}$, respectively. To tailor the recommendation-oriented features, we adopt two non-linear transformations to cast $\bm{F_{t}}$ and $\bm{F_{v}}$ into the same feature space:
(1) $\bm{q_{t}}$ $=\sigma(\textbf{W}_{t}\bm{F_{t}}+b_{t}),$
$\bm{q_{v}}$ $=\sigma(\textbf{W}_{v}\bm{F_{v}}+b_{v}),$
where $\textbf{W}_{t}$, $\textbf{W}_{v}\in\mathbb{R}^{d\times d_{0}}$ and $b_{t}$, $b_{v}\in\mathbb{R}^{d}$ denote the weight matrices and bias vectors for textual and visual modality, respectively. $\sigma(\cdot)$ is the activation function.
Following the disentangled representation learning process in DMRL (Liu et al., 2022), we first split the feature vector into $K$ chunks. Each chunk corresponds to a specific item attribute, such as price, brand, etc. For simplicity, we equally split the representation of each modality into $K$ continuous chunks. Take item ID embedding as an example:
(2) $\bm{q_{i}}=(\bm{q_{i}^{1}},\bm{q_{i}^{2}},\cdots,\bm{q_{i}^{K}}),$
where $\bm{q_{i}^{k}}\in\mathbb{R}^{\frac{d}{K}}$ is the item ID corresponding to the $k$-th attribute. Analogously, $\bm{q_{t}}=(\bm{q_{t}^{1}},\bm{q_{t}^{2}},\cdots,\bm{q_{t}^{K}})$, $\bm{q_{v}}=(\bm{q_{v}^{1}},\bm{q_{v}^{2}},\cdots,\bm{q_{v}^{K}})$ and $\bm{p_{u}}=(\bm{p_{u}^{1}},\bm{p_{u}^{2}},\cdots,\bm{p_{u}^{K}})$ are defined as the embeddings of textual feature, visual feature and user ID embedding, respectively.
3.1.3. Intuition of Attribute-driven Disentanglement
Our work not only aims to recommend accurate items to users but also advocates for attribute-driven disentanglement in multimodal recommender systems to enhance the interpretability and controllability of recommendations. Specifically, unlike the existing method that disentangles the latent factors in an unsupervised manner, we leverage the semantic labels of item attributes to learn attribute-specific sub-spaces for each attribute to obtain disentangled representations. In this paper, the vector of each modality is composed of $K$ chunks, with the assumption that each chunk is associated with an attribute (such as price or brand). The achievement of attribute-driven disentangling requires two necessary conditions. Firstly, each chunk should have a clear attribute reference, and chunks associated with different attributes should be distinguishable. For example, chunk $i$ is associated with the price, while chunk $j$ represents the vector of the brand. Secondly, each chunk of a specific attribute should accurately capture the specific attribute value. For example, chunk $i$ should reflect the price level (i.e., expensive or cheap) of the product.
3.2. Attribute-driven Disentangled Representation Learning
In this section, we elaborate on our proposed model, termed AD-DRL, an acronym for Attribute-Driven Disentangled Representation Learning model. We aim to improve the interpretability and controllability of recommendation models by assigning a specific attribute to each factor in multimodal features. Specifically, AD-DRL comprises two disentangling modules: high-level and low-level attribute-driven disentangled representation learning. In the high-level attribute-driven disentangled representation learning module, we exploit the difference between attribute factors within the same modality feature and the consistency of the same attribute factor across different modalities. In contrast, the low-level attribute-driven disentangled representation learning module leverages the intrinsic relationships between items sharing the same attribute value.
3.2.1. High-Level Attribute-driven Disentangled Representation Learning
To obtain robust and independent representations for each attribute factor in multimodal features, AD-DRL enables disentangling within each modality feature and ensures consistency of representations across modalities. Next, we detail the intra-modality disentanglement and inter-modality disentanglement.
Intra-Modality Disentanglement. After extracting the feature vector from each modality, we simply mechanically split the vector into several chunks. Nevertheless, different attribute factors are still entangled within the chunks. That is, for a chunk, it may contain both price- and brand-related features. Therefore, in order to disentangle attribute factors in each modality feature, we employ an attribute classifier to force each chunk to be sufficient to predict the corresponding attribute, as shown in Fig. 1(a). Taking the $k$-th chunk $\bm{q_{t}^{k}}$ of item textual embedding as an example,
(3) $\begin{cases}\mathbf{z}^{k}&=\mathbf{W}_{intra,t}\bm{q_{t}^{k}}+\mathbf{b}_{intra,t},\\ l_{intra,t}^{k}&=-\sum_{n=1}^{K}\tilde{z}_{n}^{k}\log\frac{\exp{z_{n}^{k}}}{\sum\nolimits_{m}\exp{z_{m}^{k}}}\end{cases}$
where $\mathbf{W}_{intra,t}\in\mathbb{R}^{K\times\frac{d}{K}}$ and $\mathbf{b}_{intra,t}\in\mathbb{R}^{K}$ are the weight matrix and bias vector of the classifier, respectively. $\mathbf{z}^{k}$ is the predicted logits and $z_{n}^{k}\in\mathbf{z}^{k}$. $\tilde{z}_{n}^{k}$ denotes the ground truth (attribute label) of chunk $\bm{q_{t}^{k}}$. For all the chunks in textual embedding $\bm{q_{t}}$, we have the following loss for all attribute factors in textual features:
(4) $l_{intra,t}=\sum_{k=1}^{K}l_{intra,t}^{k}.$
Similarly, we can define the losses $l_{intra,u}$, $l_{intra,i}$ and $l_{intra,v}$ to encourage the features of each attribute factor to be concentrated in the corresponding chunk of user ID, item ID and visual feature, respectively. Finally, the total loss for the intra-modality disentanglement is formulated as:
(5) $l_{intra}=l_{intra,u}+l_{intra,i}+l_{intra,t}+l_{intra,v}.$
Inter-Modality Disentanglement. In addition to separately disentangling each modality, a serious challenge in disentangling representation learning for multi-modal features is handling the inter-relationship between the factors disentangled from multiple modalities. Intuitively, the chunks of the same attribute factor from different modality features should be consistent. For example, for the brand attribute, an item should have the same brand information in both the visual and textual features.
Inspired by this, we disentangle attribute factors in different modality features from the perspectives of consistency in multi-modality recommendations. Two chunks of the same attribute factor from two different modality features should have a high similarity because they are related to the same attribute. In contrast, two chunks from different modality features of different attribute factors should have low similarity as they provide different attribute information.
To this end, we design a cross-modal contrastive loss as shown in Fig. 1(b). Specifically, for any two modality representations, such as $\bm{q_{t}}$ and $\bm{q_{v}}$, we take two chunks of the same attribute factor (i.e., $(\bm{q_{t}^{k}},\bm{q_{v}^{k}})$) as positive pairs, and two chunks corresponding to different attributes (i.e., $(\bm{q_{t}^{k}},\bm{q_{v}^{n\neq k}})$, $(\bm{q_{t}^{n\neq k}},\bm{q_{v}^{k}})$) as negative pairs. After that, the cross-modal contrastive loss between textual and visual features is defined as follows:
(6) $\begin{cases}l^{k}_{t\rightarrow v}&=-\log\frac{\exp((\bm{q_{t}^{k}}\cdot\bm{q_{v}^{k}})/\tau)}{\sum_{n=1}^{K}\exp((\bm{q_{t}^{k}}\cdot\bm{q_{v}^{n}})/\tau)},\\ l^{k}_{v\rightarrow t}&=-\log\frac{\exp((\bm{q_{v}}^{k}\cdot\bm{q_{t}^{k}})/\tau)}{\sum_{n=1}^{K}\exp((\bm{q_{v}^{k}}\cdot\bm{q_{t}^{n}})/\tau)},\\ l_{t\leftrightarrow v}&=\sum_{k=1}^{K}l^{k}_{t\rightarrow v}+\sum_{k=1}^{K}l^{k}_{v\rightarrow t},\end{cases}$
where $\cdot$ represents the dot product and $\tau\in\mathbb{R}^{+}$ is a scalar temperature parameter. By applying this contrastive learning constraint to any two out of the three modalities, we can achieve disentanglement and alignment across features of various modalities:
(7) $l_{inter}=l_{i\leftrightarrow t}+l_{i\leftrightarrow v}+l_{t\leftrightarrow v}.$
3.2.2. Low-level Attribute-driven Disentangled Representation Learning.
Since each factor has a clear attribute meaning, each attribute should have its corresponding specific attribute values. Specifically, each attribute $k$ is associated with a list of possible attribute values $(\tilde{y}^{k}_{1},\tilde{y}^{k}_{2},\cdots,\tilde{y}^{k}_{A_{k}})$, where $A_{K}$ is the total number of possible values for that attribute. For example, the attribute value of popularity can be stratified into five distinct levels: Super Popular, Popular, Moderate, Emerging, and Unknown. We hope that the representations disentangled based on attributes can reflect the specific attribute values of the item.
It is obvious that any single modality may not contain sufficient information for a particular attribute. For example, images may more intuitively reflect attributes such as color and brand of an item, but may not directly reflect attributes such as price or popularity. Therefore, in order to comprehensively and accurately depict the attribute values of the item, we first integrate the features from all modalities together. To achieve this, we apply a multimodal attention mechanism to measure the emphasis of different modalities on different attributes. Specifically, for the $k$-th attribute, the specific attention values to different modalities are computed by a two-layer neural network:
(8) $\begin{cases}\hat{\mathbf{a}}^{k}=\textbf{W}_{a2}\tanh(\mathbf{W}_{a1}(\bm{q_{i}^{k}}+\bm{q_{t}^{k}}+\bm{q_{v}^{k}})+\bm{b}_{a}^{k}),\\ \mathbf{a}^{k}=Softmax(\hat{\mathbf{a}}^{k}),\end{cases}$
where $\mathbf{W}_{a1}\in\mathbb{R}^{3\times\frac{d}{K}}$ and $\textbf{W}_{a2}\in\mathbb{R}^{3\times 3}$ are the weight matrices corresponding to the first and second layers of the neural network, respectively. $\bm{b}_{a}^{k}$ denotes the bias vector and $\tanh$ is the activation function. $Softmax$ is adopted to normalize $\hat{\mathbf{a}}^{k}$ to a probability distribution.
Thereafter, we obtain the final representation of the $k$-th attribute factor for the item as follows:
(9) $\bm{q_{y}^{k}}=a_{i}^{k}\cdot\bm{q_{i}^{k}}+a_{t}^{k}\cdot\bm{q_{t}^{k}}+a_{v}^{k}\cdot\bm{q_{v}^{k}},$
where $a_{i}^{k}$, $a_{t}^{k}$, $a_{v}^{k}\in\mathbf{a}^{k}$ are the attention weights of item ID embedding, textual feature and visual feature, respectively.
Similar to the disentangling at the intra-modality disentanglement in Section 3.2.1, we aim to achieve disentangled representation learning at the low level through attribute value prediction via a classification layer (as shown in Fig. 1(c)):
(10) $\begin{cases}\mathbf{y}^{k}&=\mathbf{W}_{k}\bm{q_{y}^{k}}+\mathbf{b}_{k},\\ l^{k}_{low}&=-\sum_{i=1}^{A_{K}}\tilde{y}_{i}^{k}\log\frac{\exp{y_{i}^{k}}}{\sum\nolimits_{j}\exp{y_{j}^{k}}},\\ l_{low}&=\sum_{k=1}^{K}l^{k}_{low}\end{cases}$
where $\mathbf{W_{k}}\in\mathbb{R}^{K\times\frac{d}{K}}$ and $\mathbf{b}_{k}\in\mathbb{R}^{K}$ are the weight matrices and bias vector of the classifier, respectively. We supervise the training of each attribute subspace via independent attribute classification tasks defined in the form of a cross-entropy loss.
3.3. Preference Prediction and Model Learning
3.3.1. Preference Prediction
So far we have discussed how to obtain attribute-driven disentangled representations $\bm{p_{u}}=(\bm{p_{u}}^{1},\bm{p_{u}}^{2},\cdots,\bm{p_{u}}^{K})$ and $\bm{q_{y}}=(\bm{q_{y}^{1}},\bm{q_{y}^{2}},\cdots,\bm{q_{y}^{K}})$ for the user $u$ and item $i$, respectively. By assigning each disentangled representation with an attribute as described above, we are able to predict users’ preferences at the attribute level. This enhances the interpretability and controllability of our model. Specifically, to estimate a user $u$’s preference for an item $i$, it is crucial to consider her preference for each attribute of the item. To achieve this, we first compute the user’s preference score for each attribute of the item, and subsequently, integrate the scores of each attribute to estimate the user’s overall preference for the item:
(11) $\begin{cases}s_{u,i,k}=\sigma(\bm{p_{u}^{k}}\cdot\bm{q_{y}^{k}}),\\ s_{u,i}=\sum_{k=1}^{K}s_{u,i,k},\end{cases}$
where $\cdot$ symbolizes the dot product and $\sigma(\cdot)$ denotes the activation function. In this paper, we use a softplus function to ensure the resultant score $s_{u,i,k}$ is positive. $s_{u,i,k}$ denotes the user $u$’s preference score for the attribute $k$ of the item $i$.
3.3.2. Training Protocol
Based on the predicted preference score above, we recommend a set of top-$n$ ranked items which match the target user’s preferences. The Bayesian Personalized Ranking (BPR) loss function (Rendle et al., 2009) is employed to optimize the model parameters $\bm{\Theta}$,
(12) $\mathcal{L}_{BPR}=\sum_{(u,i_{+},i_{-})\in\mathcal{D}}-\log\phi(s_{u,i_{+}}-s_{u,i_{-}})+\lambda{\lVert\bm{\Theta}\rVert}_{2}^{2},$
where $\lambda$ is the coefficient controlling $L_{2}$ regularization; $\mathcal{D}$ denotes the training set; $i_{+}$ and $i_{-}$ are the observed and unobserved item in the interaction records of user $u$, respectively. Overall, the total loss of AD-DRL is formulated as,
(13) $\mathcal{L}=\mathcal{L}_{BPR}+\alpha{\sum_{(u,i)\in\mathcal{D}}l_{intra}}+\beta{\sum_{(u,i)\in\mathcal{D}}l_{inter}}+\gamma{\sum_{(u,i)\in\mathcal{D}}l_{low}},$
where $\alpha$, $\beta$ and $\gamma$ are the hyperparameters to control the weight of three disentanglement module.
4. Experiments
Table 1. Basic statistics of the three datasets. ”#” denotes the number of statistical values.
Dataset #user #item #interaction sparsity #price #popularity #brand #category
Baby 12,637 18,646 121,651 99.95% 5 5 663 1
Toys Games 18,748 30,420 161,653 99.97% 5 5 1,288 19
Sports 21,400 36,224 209,944 99.97% 5 5 2,081 18
We conducted extensive experiments on three benchmark datasets to validate the effectiveness of the proposed method. In particular, the experiments are mainly performed to address the following research questions:
  • RQ1: Can the proposed disentangled representation learning method outperform existing state-of-the-art models?
  • RQ2: Can AD-DRL effectively learn disentangled representations by utilizing attributes?
  • RQ3: How do the controllability and interpretability of our proposed AD-DRL perform?
  • RQ4: How do the key components of AD-DRL, specifically high- and low-level attribute-driven disentangled representation learning, contribute to the overall performance of AD-DRL on different datasets?
4.1. Experimental Setup
In the following, we first provide the basic information of the evaluated benchmark datasets. We then detail the standard evaluation metric, followed by the key baselines used in this work, and we end this section with some implementation details.
4.1.1. Datasets
We use the widely adopted real-world recommendation dataset, the Amazon review dataset [^1] http://jmcauley.ucsd.edu/data/amazon. (McAuley and Leskovec, 2013), for evaluation in our experiments. Apart from user-item interaction data, this dataset also includes multimodal information (i.e., reviews and images) and various attributes (i.e., price, brand, etc.) of the items on 24 product categories. Three product categories from this dataset are used in the evaluation: Baby, Toys Games and Sports. Following (Liu et al., 2022), for all datasets, unpopular items and inactive users are filtered out to ensure that all the items and users have at least 5 interaction records. The basic statistics of the three datasets are shown in Table 1.
In our setting, in addition to the user-item interaction data and multimodal information of items, multiple attributes and their associated attribute values of items are required to guide the disentangled representation learning process. Specifically, we adopt four typical attributes (i.e., price, popularity  [^2] Popularity is calculated by counting the number of times each item is purchased., brand and category), and the attribute values for each attribute are compiled based on the metadata provided by the Amazon dataset: for brand and category values, we directly used the values provided by Amazon; for price and popularity, following the methods used by CoHHN (Zhang et al., 2022), we discretize them into five levels according to their numerical values. Table 1 shows the specific statistical information for each attribute in our experiments.
4.1.2. Baselines
We compared our method with the state-of-the-art methods, including both the CF-based methods (i.e., NeuMF(He et al., 2017), NGCF (Wang et al., 2019) and DGCF (Wang et al., 2020)) and Multimodal CF-based methods (JRL (Zhang et al., 2017b), MMGCN (Wei et al., 2019), MAML (Liu et al., 2019), GRCN (Wei et al., 2020), BM3 (Zhou et al., 2023) and DMRL (Liu et al., 2022)). For each type of method, we also select a disentangled representation learning model (i.e., DGCF and DMRL) for comparison.
4.1.3. Evaluation Metrics and Parameter Settings
For each dataset, we randomly split the interactions from each user with an $8:2$ ratio to construct training and testing sets. We randomly selected 10% of interactions from the training set as a validation set for hyper-parameter tuning. To evaluate the performance, we focus on the top-$n$ recommendation task, aiming to return the top-$n$ items as recommendations. We adopt two widely used evaluation metrics, i.e., Recall@$n$ and NDCG@$n$, to evaluate the accuracy, where $n$ is set as 20 by default.
The Pytorch toolkit (Paszke et al., 2019) is utilized to implement our models. To ensure a fair comparison, we optimize all the methods with Adam optimizer (Kingma and Ba, 2015), where the default learning rate and batch size are set to 0.0001 and 1024, respectively. We fixed the embedding size of each factor to 32 for AD-DRL and its variants on all datasets. The Xavier initializer (Glorot and Bengio, 2010) is used to initialize the model parameters. The $\alpha$, $\beta$, $\gamma$ and $L_{2}$ regularization coefficient are searched in $\{1e^{-5},5e^{-5},1e^{-4},\cdots,5e^{+0},1e^{+1}\}$. The number of negative examples is searched in $\{2,4,8\}$. Besides, model parameters are saved every five epochs. Early stopping strategy (Wang et al., 2019) is performed, i.e., premature stopping if Recall@20 does not increase for 50 successive epochs.
4.2. Performance Comparison (RQ1)
Table 2. Performance comparison of different recommendation methods over three datasets. The best results are highlighted in bold.
Datasets Baby Toys Games Sports
Metrics Recall NDCG Recall NDCG Recall NDCG
NeuMF 0.0502 0.0224 0.0253 0.0128 0.0330 0.0157
NGCF 0.0694 0.0313 0.0970 0.0587 0.0707 0.0337
DGCF 0.0788 0.0465 0.1262 0.1085 0.1026 0.0629
JRL 0.0579 0.0266 0.0472 0.0413 0.0368 0.0214
MMGCN 0.0814 0.0496 0.1171 0.1065 0.0913 0.0572
MAML 0.0867 0.0521 0.1183 0.1117 0.1029 0.0676
GRCN 0.0883 0.0541 0.1336 0.1236 0.1065 0.0693
DMRL 0.0906 0.0561 0.1434 0.1331 0.1111 0.0711
BM3 0.0911 0.0424 0.1147 0.0683 0.1121 0.0536
AD-DRL 0.0968* 0.0588* 0.1524* 0.1435* 0.1200* 0.0756*
Improv. 6.84% 4.81% 6.28% 7.81% 8.01% 6.33%
  • The symbol * denotes that the improvement is significant with $p-value<0.05$ based on a two-tailed paired t-test.
Figure 2. Visualization of entangled and disentangled vectors from different modalities. For the second row, different colors represent different attributes. Specifically, green represents price, purple represents popularity, brown represents brand, yellow represents category, and blue represents others. For the first row, since the model does not involve attributes, different colors only represent different factors for ease of distinction.
We summarize the overall performance comparison results in Table 2. The methods in the first block solely use the user-item interactions, while the methods in the second block also exploit textual and visual information along with the user-item interactions. From this table, we can have the following observations:
  • NeuMF, NGCF, and DGCF are the methods that are solely trained using user-item interactions. Among them, NeuMF can achieve promising performance compared to conventional MF-based methods (He et al., 2017). This is attributed to the capacity of deep neural networks to model the non-linear interactions between users and items. Benefiting from the high-order information, both NGCF and DGCF accomplish state-of-the-art results. Furthermore, DGCF performs better than NGCF by characterizing diverse user intents through the disentangled representation technique. This highlights the effectiveness of disentangled representation learning, particularly in improving the robustness of user and item representations.
  • In general, the multimodal recommendation methods perform better than those only using user-item interactions, demonstrating the effectiveness of leveraging multimodal information on learning user and item representations. Although the simple neural network structure results in lower performance of JRL compared to the graph-based methods (NDCG and DGCF), incorporating rich multi-modal information enables it to outperform NeuMF. By exploiting user-item interactions to guide the representation learning in different modalities, MMGCN yields better performance over NGCF on all the datasets. MAML outperforms NGCF by modeling users’ diverse preferences by using multimodal features of items. GRCN surpasses MAML by employing modality features to discover and prune potential false-positive edges on the user-item interaction graph. DMRL performs better than the other baselines over almost all the datasets. It should be credited to its capability to capture the different contributions of features from different modalities for each disentangled factor.
  • Furthermore, our proposed AD-DRL model consistently outperforms all baseline methods over all three datasets by a large margin. We credit this to the joint effects of the following two aspects. Firstly, robust user and item representations can be derived using attributes at different levels to guide the disentangled representation learning process. Secondly, incorporating the additional attribute information into representation learning can alleviate the data sparsity problem in recommendation.
4.3. Visualization of Disentangled Representations (RQ2)
[…]
Figure 3. The preference scores of two users ($u$18629 and $u$6805) for different attributes of two items ($i$2099 and $i$3460).
4.4. Interpretability Study (RQ3)
To gain a deeper insight into the interpretability of our model, in this section, we provide some qualitative examples in Figure 3. As an illustrative example, we randomly sample two users (u18629 and u6805) from the Sports dataset who purchased the same items (i2099 and i3460). Since we have disentangled the representation of users and items based on the attributes, we can analyze how each attribute contributes to recommendation results based on Equation 11. This greatly enhances the interpretability of our model. For example, price contributes most to the interaction (u10826, i2099). This suggests that u10826 prefer i2099 due to its price. However, u6805 is more likely to purchase the i2099 because of her preference for its brand. Such observation indicates that different users have diverse preferences for the same product. In addition, we can also see that user preferences exhibit consistency. For example, u6905 values the brand of the product more than its popularity or price, which is different from u10826.
4.5. Controllability Study (RQ3)
One of the primary targets of our method is to improve the controllability of recommendation systems. In this section, we assess the controllability of AD-DRL by removing the representation chunks of factors associated with price and popularity. As shown in Figure 4(a), for the attribute of price, we selected 100 users who always purchase inexpensive products based on their purchase records. ”w/ price” indicates the recommendation results of the AD-DRL model utilizing all factors, where the model tends to recommend inexpensive products to users. However, when we removed the chunk associated with the price (denoted as ”w/o price” in Figure 4(a)), we found that the recommended products from AD-DRL were more diverse in terms of price. This underscores AD-DRL’s capability for attribute-level controllability. It is with noting that after removing the price attribute, the products recommended by AD-DRL still tend to be inexpensive. This could be attributed to the retained chunk associated with the brand, and there are correlations between brand and price. For example, brand reputation plays a significant role in pricing. Some high-end brands offer products at a premium price, while others provide more affordable options. Similar results can be seen in another group of users who consistently purchase expensive products. As a comparison, due to the lack of a clear definition for each factor in previous work, the results of removing the factors associated with specific attributes are unpredictable.
Figure 4. Statistics of the recommendation results while AD-DRL removes the chunk of price factor or popularity factor. ##\##counts refer to the number of items with different attribute values.
4.6. Ablation Study (RQ4)
Table 3. Ablation study of our proposed AD-DRL method over three datasets. The best results are highlighted in bold.
Datasets Baby Toys Games Sports
Metrics Recall NDCG Recall NDCG Recall NDCG
[…]
AD-DRL 0.0968 0.0588 0.1524 0.1435 0.1200 0.0756
[…]
5. Conclusion
Existing disentangled representation learning methods in recommender systems disentangle the underlying factors behind user-item interaction data in an unsupervised manner, leading to limit interpretability and controllability of recommendation. This paper proposes an attribute-driven disentangled representation learning method, termed AD-DRL, which disentangles attribute factors in various multimodal features. For obtaining robust and independent representations, the proposed AD-DRL enables disentangling within each modality feature and ensures consistency of representations across modalities at the attribute level. Moreover, it leverages the intrinsic relationships between items sharing the same attribute value. Experimental results demonstrate the superiority of our method AD-DRL compared to existing methods and showcase its capability in terms of interpretability and controllability.